OpenStudy (anonymous):

Solve the equation. Show all work. All final answers must be in simplest form. 1/4+b=10

jimthompson5910 (jim_thompson5910):

1/4+b=10 b=10-1/4 b=40/4-1/4 b=39/4

OpenStudy (anonymous):

b =10 - 1/4 40/4 - 1/4 = 39/4 Jim cheats